OpenStudy (anonymous):

Two lovers have a spat and swear they will never see each other again. The girl walks due south at 6 mph while the boy walks 10 mph on a heading of North 60. Degrees West. How fast is the distance between them changing 30 minutes later?

OpenStudy (aum):



OpenStudy (aum):

After "t" hours, OG = 6*t miles OB = 10*t miles. Find GB using the law of cosines. Then find the derivative with respect to t to find the rate of change. Put t = 1/2 to find the rate of change half hour later.

OpenStudy (anonymous):

What formula do you use to find the derivative? I found GB and understand the last step.

OpenStudy (anonymous):

Into the derivative of the law of cos?